Weights of the Pacific yellowfin tuna follow a normal distribution with mean 68 pounds and standard deviation 12 pounds. For randomly caught Pacific yellowfin tuna, what is the probability that the weight is less than 50 lbs?
Answer by ewatrrr(24785)

 
Hi
mean 68 pounds and standard deviation 12 pounds.
P(≤50) = P(z≤(50-68)/12) = P(z ≤ -1.5) = .0668